OpenStudy (anonymous):

A doctor has ordered that a patient be given 15g of glucose, which is available in a concentration of 75.00g glucose/1000.0 mL of solution.What volume of solution should be given to the patient?

OpenStudy (anonymous):

200ml

OpenStudy (aaronq):

lol at least show him how you did it

OpenStudy (anonymous):

in 1000ml of solution 75gm of glucose is present we need only 15gm i.e15/75=1/5th of total 1000ml =1/5*1000=200ml

OpenStudy (aaronq):

nice (Y) or you can use the density formula: \(\rho=\dfrac{m}{V}\)